The fan cost index tracks the average costs for attending sporting events, including tickets, drinks, food, parking, programs, and souvenirs. According to the FCI, a family of four would spend a total of $592.30 to attend two Major League Baseball (MLB) games and one National Basketball Association (NBA) game. The family would spend $691.31 to attend one MLB and two NBA games. Write and Solve a system of equations to find the family's costs for each kind of games according to the FCI.
Answer by stanbon(75887)

According to the FCI, a family of four would spend a total of $592.30 to attend two Major League Baseball (MLB) games and one National Basketball Association (NBA) game. The family would spend $691.31 to attend one MLB and two NBA games. Write and Solve a system of equations to find the family's costs for each kind of games according to the FCI.
-----
Equation:
2B + N = 592.30
B +2N = 691.31
-----------------------
Multiply thru the 1st equation:
4B + 2N = 2*592.30
B + 2N = 691.31
--------------------------
Subtract and solve for "B"
3B = $493.29
B = $164.43
==================
Solve for "N":
2B + N = 592.30
328.86 + N = 592.30
N = $263.44
===================
Cheers,
Stan H.
===================